EGS Abstract for Nice, 2000

EARTHQUAKE TRIGGERING BY STATIC AND DYNAMIC STRESS CHANGES

M.E. Belardinelli (1), A. Bizzarri (2) and M. Cocco (2)

(1) Settore di Geofisica, University of Bologna, Viale Berti-Pichat 8, 40127, Bologna, Italy, (2) Istituto Nazionale di Geofisica, Via di Vigna Murata 605, Roma, Italy.

elina@ibogfs.df.unibo.it

Different theoretical modeling and observational tests have shown that earthquakes perturb the state of stress of a region, possibly encouraging the subsequent failure of neighboring faults. The stress perturbation due to an earthquake can be represented as a step function (the static stress change) with a transient signal superimposed, here referred as dynamic stress change. Induced failures are generally delayed with respect to the application of the coseismic stress perturbation due to a previous event occurred nearby. Induced failures can be also advanced with respect to the failure time (unknown) expected in unperturbed conditions. We use a spring-slider model with a rateand state- dependent friction law to investigate the effect of dynamic and static stress changes on earthquake triggering. The effects of the static and dynamic imposed traction are studied separately using analytical expressions. We also consider the response to a traction time history inferred by direct modeling of an extended seismic source. Our preliminary results show that the join action of the static and the dynamic stress changes is likely to promote subsequent failures.

Abstracts to be submitted on or before December 15, 1999 to

EGS Office Max-Planck-Str. 13 37191 Katlenburg-Lindau Germany Tel.: [+49] 5556-1440 Fax.: [+49] 5556-4709

                                                    Email: EGS@Copernicus.org
                                      http://www.copernicus.org/EGS/EGS.html

Submittal Information

1. Conference EGS, 2000

2. Submission type First submission 3.

Title

Earthquake triggering by static and dynamic stress changes

4.

Author(s)

BELARDINELLI, M.E.; BIZZARRI, A.; COCCO, M.

5.

Session SE22 04 Earthquakes. Earthquake fault interaction through stress transfer

6. Organizer  Dr. M. Cocco
7. Equipment  NONE

8. Support Award

NONE

9. Presentation Oral presentation strongly preferred 10. Abstract type

L^{A}TEX

11. Contact author Belardinelli

12. First name  Maria Elina
13. Title  Dr.

14.

Department

Settore di Geofisica of the Physics Department

15. Organization University of Bologna 16.

Address

Viale Berti-Pichat 8

17. Postal code  40127
18. City  Bologna
19. Country  Italy
20. Telephone  +39 051 2095018
21. Telefax  +39 051 2095058

22.

Email

                       elina@ibogfs.df.unibo.it